OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a base1 = 13 and a base17 = –51 ? and can someone tell me how to make it a base number instead of typing ''base''

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

I am not sure about the answer nyan but @lgbasallote will surely help you in this!! hope he helps!

OpenStudy (lgbasallote):

you right a_{base} in your equations editor i.e. \(a_{1}\) is a_{1}

OpenStudy (lgbasallote):

you write* lol

OpenStudy (anonymous):

a_1?

OpenStudy (lgbasallote):

anyway the general formula is \(\Large a_{n} = a_{1} + (n - 1)d\) first thing you need to do is to find d... you're going to perform arithmetic sequence twice...the firs is \(\large a_{17} = 13 + (n-1)d\) \(\large -51 = 13 + (17-1)d\) \(\large -51 - 13 = 16d\) \(\large -64 = 16d\) d = -4 with that given info solve for \(a_{35}\) hint: a_1 = 13 n = 35 d = -4

OpenStudy (anonymous):

i dont quite understand the formula but to be honest i was put in an online algebra 2 class thats way too advanced for me i have ADHD and dyslexia, i will give medles for the answers, the sooner i finish this the sooner you guys can get rid of me lol :)

OpenStudy (anonymous):

also i dont have a calculator

OpenStudy (lgbasallote):

hmm one...i dont really care bout the medals :) i just help people ...two...you don't need a calculator for this...three....im sure this is adequate for ADHD and dyslexia people dont worry ^_^

OpenStudy (anonymous):

but i still dont understand what im suposta do?

OpenStudy (anonymous):

its one of these –126 –123 –120 –117